OpenStudy (anonymous):

hello guys, can someone give me an explanation and a couple of exercises on how to solve some basic differential equations? Thank you beforehand

OpenStudy (anonymous):

thank you

OpenStudy (tanya123):

Did you get what you were finding ?

OpenStudy (anonymous):

yup, thank you very much

OpenStudy (tanya123):

All right! your welcome =')

OpenStudy (theeric):

http://tutorial.math.lamar.edu/Classes/DE/DE.aspx Those are dry, but can be helpful! :)

OpenStudy (unklerhaukus):

Solve the differential equation for \(y(x)\), given the initial condition \(y_0=y(0)\)\[\frac{\mathrm dy}{\mathrm dx}=x,\qquad\qquad y_0=1\]__________________________________ Separate the variables, and integrate \[\begin{align}\mathrm dy& =x\,\mathrm dx\\\int\mathrm dy&=\int x\mathrm dx\\y&=x^2/2+c\end{align}\] Applying the initial condition, we find the constant of integration \[\qquad y(0)=c=1\] The solution is : \[y(x)=x^2/2+1\] Differentiating the solution gives back the original DE as expected. \[\frac{\mathrm dy}{\mathrm dx}=x\]